OpenStudy (anonymous):

A lottery winner must decide between two methods of payment. Choice one is to receive a lump sum of $100,000,000. Choice two is to begin with one cent on day 1, five cents on day 2, 25 cents on day three, and so on until the end of 16 days. Which choice should the lottery winner select, how do you solve these??? Can someone please break it down and explain??? pleasee

OpenStudy (anonymous):

.01+.05 + .25 if it is geomtric series whichh i suspect from your other qs then .01*r^2=.25 SFR r=5 So .01*(5)^16=1525878906.00

OpenStudy (anonymous):

100,000,000-1525878906 100000000 1525878906 is clearly more :D

OpenStudy (precal):

choice 2, this is a classic math problem king gives this servant this choice for saving his daughter (I think), servant picks choice 2 and king goes broke

OpenStudy (anonymous):

$1425878906 more :D

OpenStudy (anonymous):

okay thank you again! your theee best!!(:

OpenStudy (anonymous):

aww *blush* are not :)

OpenStudy (precal):

go with timo86m solution, it is more mathematical, mine is just historical or urban legend like

OpenStudy (anonymous):

haha yesss huh!!

OpenStudy (anonymous):

there is another story of the inventor of chess. the king liked it so much he said what do you want for it. he said one grain of wheat for square 1 then 2 for square 2 and 4 for square 3 and keep doubling it :) each square I think the king accepted and went broke.

OpenStudy (precal):

I think it is the same story.........

OpenStudy (precal):

see I bet that asker learned more about this problem than they thought they would get

OpenStudy (anonymous):

A ball is dropped from the top of a 1,000 foot building. The height of the ball is half its original height after each bounce. What will the height of the ball be after 10 bounces? how would I SOLVE this

OpenStudy (anonymous):

1000*.5^10 I think not sure tho

OpenStudy (anonymous):

1000/2^10

OpenStudy (precal):

geometric sequence by after 10 bounces, I think you power should be 11 -that accounts for the 10 bounce to be completed

OpenStudy (anonymous):

nah i sure it is 10

OpenStudy (anonymous):

1000/2^1= 500 1000/2^2=250 3 4 5 6 7 8 9 1000/2^10 = 125/128

OpenStudy (anonymous):

0 counts as your 11th term right?

OpenStudy (precal):

I just remember these word problems were always the power +1, I always had to draw them out

OpenStudy (anonymous):

it is confusing after jsut doing geometric series problems lol. But ^0 also counts so it is ^10 not ^11 right? but idk you the one with higher smart score tho :)

OpenStudy (precal):

yes but I haven't done these problems in a while, I tended to skip them because I always forgot if it is 10 or 11 power, hope their teacher believes in partial credit

OpenStudy (anonymous):

lol if it is an online class no such luck :P Oh and I haven't done them in a while either :P BUt drawing it out helps

OpenStudy (precal):

see another reason a real live teacher is better, you can negotiate with a human and not a machine

OpenStudy (anonymous):

Threaten the machine with a virus or and EMP :P lol

OpenStudy (precal):

yes of course